A researcher wishes to determine the proportion of college students who smoke. He determined that from a sample of 150 college students, 35admitted that they smoke.Is it valid to infer that 23% of college students smoke? Use a = 0.05

1
Expert's answer
2022-04-29T05:31:11-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.23"

"H_a:p\not=0.23"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{\dfrac{35}{150}-0.23}{\sqrt{\dfrac{0.23(1-0.23)}{150}}}\approx0.0970"

The p-value is "p =2P(Z>0.0970)= 0.922726"

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

Since it is observed that "z = 0.0970 \le1.96= z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p =0.922726," and since "p = 0.922726 \ge 0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than 0.23, at the "\alpha = 0.05" significance level.